Give an example that the commutative law does not holds on the matrix multiplication.


1
Expert's answer
2021-10-11T15:55:39-0400

Since


"\begin{pmatrix}
2 & -2\\
0 & 1
\end{pmatrix}
\begin{pmatrix}
-1 & 2\\
0 & 1
\end{pmatrix}
=
\begin{pmatrix}
-2 & 2\\
0 & 1
\end{pmatrix}," but


"\begin{pmatrix}
-1 & 2\\
0 & 1
\end{pmatrix}
\begin{pmatrix}
2 & -2\\
0 & 1
\end{pmatrix}
=
\begin{pmatrix}
-2 & 4\\
0 & 1
\end{pmatrix},"


we conclude that the commutative property does not hold on the matrix multiplication.